OpenStudy (anonymous):

solve the given formula to find specified variable. 1/s+1/q=1/h Find h.

OpenStudy (anonymous):

you can do reciprocal, but remember applied it to the right side AND the left side. \[\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }=\frac{ 1 }{ x }\]\[2+4=x\]

OpenStudy (anonymous):

how did you get 2 and 4 o: if that's the case x = 6?

OpenStudy (anonymous):